# Evaporative damping in open system theory of Bose-Einstein Condensates

### v1

@@ I Introduction
Commonly only one mechanism, so called number damping ((a) in figure 1), is used to mdel dissipation in the SPGPE[28]. It stems from an exchange of particles between the thermal cloud and the coherent region. Number damping has been succesfully applied to calculate and prepare grand-canonical equilibrium distributions [29, 30, 31, 32, 33]. However, if employed to describe dynamics, it does not match observations and is often arbitrarily enlarged [34, 35, 36, 37, 38, 39, 40, 41].
@@
Next to number and energy damping, a third process ((c) in figure 1) has been so far discarded in the derivation of the SPGPE as it was claimed to be small. It stems from an evaporation of coherent particles into the incoherent cloud (and its inverse process of a particle capture by the coherent region). We derive the effect of this evaporative damping on the c-field in the following and conclude that it is not negligible but rather comparable in strength to the number damping process. As most of the physics occurs in the $C$-region, it is driven by highly nonlinear c-field dynamics. This evaporative damping can be cast in a form (56) that consists of a contribution resembling energy damping (corresponding to the lower particle of figure 1(c)) and one that resembles number damping (corresponding to the upper particle of figure 1(c)). If we were to choose the cut-off high enough that all trapped particles are in the coherent region, the other mechanisms in the SPGPE, number and energy damping, vanish. Evaporative damping, however, remains, then describing the evaporation of particles out of the trap.
@@ II.1 SPGPE master equation
Later, we will also include the influence of the mean field shift due to the interaction energy. The SPGPE is then an evolution equation for a classical wave containing information about the low energy states.
@@ III.2 Energy scale
This is, we assume that the particle in the incoherent region participating in the scattering has typically an energy just above the cut-off, where the thermal distribution is highest. The particles in the coherent region likewise are then choosen from those states that are highest populated, while still carrying enough energy for the evaporation. It also implies that we weaken the strict energy conservation in the scattering, the implicit momentum conservation, however, remains. One might wonder if setting $\omega$ to $\omega_{\text{cut}}$ in $R^{(+)}$ is justified if we consider the exponential up to linear order (41). To see this we expand $R^{(+)}$ in lowest order around $\omega_{\text{cut}}$ (for $\hbar\omega\geq\hbar\omega_{\text{cut}}$)
@@
Since we expect $\omega-\omega_{\text{cut}}\ll\omega-\mu/\hbar$ we conclude that the approximations
@@ III.3 Phase space picture
Inserting the constant energy approximation into (42) and (43) allows us to proceed in the derivation. We first consider the diffusion (42). Performing a Wigner-Weyl transformation and keeping only terms with second order derivatives5550th and 1st order derivatives vanish. Higher order derivatives are neglected in consistency with the TWA. we have
@@ IV.2 Approximate (Pseudo-)local form
[…]
@@
Close to the critical temperature (and hence at phase space density of order one $n_{C}\lambda_{\text{th}}^{3}\sim 1$) the coherence length will be on the order of a healing length and the integral hence on the order of one, so that we have in this regime
@@ V Dimensional Reduction
We derived the evaporative damping mechanism in three dimensions and analyzed its importance. In the following we will argue analogously to the three dimensional case to derive a two- and one-dimensional reduced form of the evaporative damping. We start with equation (32). We consider the situation in which along one or two axes a deep harmonic trap is applied. We assume that all states in the coherent region are along the axes we reduce in the single particle ground state. We can thus write in the two and one dimensional case
@@
The noises fullfill the correlations
@@
We derived a dimensionally reduced version of the evaporative damping. We note, however, that in dimensionally reduced dynamics the evaporative damping might be weakened compared to the other damping mechanisms, at least as long as the thermal cloud stays three dimensional: the allowed states in the incoherent region participating in evaporative damping have to be in the ground state along the reduced dimensions. This restriction neither applies for energy nor for number damping[46].
@@ VI Conclusions
We also derive a dimensionally reduced form of evaporative damping, equation (84). The reduction is valid if a deep harmonic trapping potential is applied along one (two) of the axes. It allows the study of the effect of evaporative damping on physics in reduced dimensions. One application could be the dynamics of stable non-linear excitations of the GPE, e.g. dark solitons [3, 4, 5, 47] in one dimension and vortices [42, 8, 9, 10] and Jones-Roberts solitons [6, 43, 7, 48] in two dimensions.

### v2

@@ I Introduction
Commonly only one mechanism, so called number damping ((a) in figure 1), is used to model dissipation in the SPGPE[28]. It stems from an exchange of particles between the thermal cloud and the coherent region. Number damping has been succesfully applied to calculate and prepare grand-canonical equilibrium distributions [29, 30, 31, 32, 33]. However, if employed to describe dynamics, it does not match observations and is often arbitrarily enlarged [34, 35, 36, 37, 38, 39, 40, 41].
@@
Next to number and energy damping, a third process ((c) in figure 1) has been so far discarded in the derivation of the SPGPE as it was claimed to be small [^1] A similiar process has been investigated in [45] in the context of an explosion of a collapsing BEC. Evaporation out of the condensate was studied, while in the context of the SPGPE we study evaporation out of the coherent region, encompassing more than just the condensate. The process was considered in context of the finite-temperature GPE (FTGPE)[22]. However, to our best knwoledge it was never considered in the SPGPE framework.. It stems from an evaporation of coherent particles into the incoherent cloud (and its inverse process of a particle capture by the coherent region). In other words, it describes scattering events between two particles in the coherent region in which one of them gains enough energy to enter the incoherent region, thereby reducing the energy of the other particle. We derive the effect of this evaporative damping on the c-field in the following and conclude that it is not negligible but rather comparable in strength to the number damping process. As most of the physics occurs in the $C$-region, it is driven by highly nonlinear c-field dynamics. This evaporative damping can be cast in a form (56) that consists of a contribution resembling energy damping (corresponding to the lower particle of figure 1(c)) and one that resembles number damping (corresponding to the upper particle of figure 1(c)). If we were to choose the cut-off high enough that all trapped particles are in the coherent region, the other mechanisms in the SPGPE, number and energy damping, vanish. Evaporative damping, however, remains, then describing the evaporation of particles out of the trap.
@@ II.1 SPGPE master equation
In SPGPE treatment, the cut-off should be chosen high enough that eigenstates of the full (interacting) Hamiltonian can be approximated by the eigenstates of the single particle Hamiltonian and low enough that the truncated Wigner approximation is valid. In practice, this leads to a cut-off several times the chemical potential $\mu$[26, 44]. Later, we will also include the influence of the mean field shift due to the interaction energy. The SPGPE is then an evolution equation for a classical wave containing information about the low energy states.
@@ III.2 Energy scale
This is, we assume that the majority of particles in the incoherent region participating in the scattering have an energy just above the cut-off, where the thermal distribution is highest. The particles in the coherent region likewise are then choosen from those states that are highest populated, while still carrying enough energy for the evaporation. It also implies that we weaken the strict energy conservation in the scattering, the implicit momentum conservation, however, remains. In the following we argue that setting $\omega$ to $\omega_{\text{cut}}$ in $R^{(+)}$ is justified while we need to consider the exponential up to linear order (41). To see this we expand $R^{(+)}$ in lowest order around $\omega_{\text{cut}}$ (for $\hbar\omega\geq\hbar\omega_{\text{cut}}$)
@@
Since we expect the energies of the evaporated particles to be close to the cut-off we have $\omega-\omega_{\text{cut}}\ll\omega-\mu/\hbar$ and we conclude that the approximations
@@ III.3 Phase space picture
Inserting the constant energy approximation into (42) and (43) allows us to proceed in the derivation. We first consider the diffusion (42). Performing a Wigner-Weyl transformation and keeping only terms with second order derivatives6660th and 1st order derivatives vanish. Higher order derivatives are neglected in consistency with the TWA. we have
@@ IV.2 Approximate (Pseudo-)local form
For temperatures above the critical temperature there is only short length coherence. Hence, $\psi^{*}(\textbf{r})\psi(\textbf{r}^{\prime})$ quickly approaches zero with increasing $\text{{{v}=}}\textbf{r}-\textbf{r}^{\prime}$ except for random fluctuations and we can assume that $\sigma$ is sharp in v. Thus, in the other terms in the integrals $\textbf{r}^{\prime}\simeq\textbf{r}$ and $\text{Im}\{\sigma\}\simeq 0$. We can argue similarly below the critical temperature if the trap is approximately flat on the scale of the healing length $\xi=\hbar/\sqrt{mn_{C}g}$. Then, for $|\text{{{v}}}|\gg\xi$ but $|\text{{{v}}}|$ small compared to the trap extension, $\langle\psi^{*}(\textbf{r})\psi(\text{{$\textbf{r}^{\prime}$}})\rangle$ is a constant and considering that the integral over the sinc gives zero we can replace $\psi^{*}(\textbf{r})\psi(\textbf{r}^{\prime})\rightarrow\psi^{*}(\textbf{r})\psi(\textbf{r}^{\prime})-\langle\psi^{*}(\textbf{r})\psi(\text{{$\textbf{r}^{\prime}$}})\rangle$. We are then again left with only fluctuations for $|\textbf{r}-\textbf{r}^{\prime}|\gg\xi$. Thus, the approximations are possible at all temperatures. We define
@@
Close to the critical temperature (and hence at phase space density of order one $n_{C}\lambda_{\text{th}}^{3}\sim 1$ [^7] This assumption requires that the coherent density is still comparable to the total density, which in the SPGPE close to the transition point is typically true.) the coherence length will be on the order of a healing length and the integral hence on the order of one, so that we have in this regime
@@ V Dimensional Reduction
We derived the evaporative damping mechanism in three dimensions and analyzed its importance. In the following we will argue analogously to the three dimensional case to derive a two- and one-dimensional reduced form of the evaporative damping. We start with equation (32). We consider the situation in which along one or two axes a deep harmonic trap is applied. We assume that all states in the coherent region are along the axes we reduce in the single particle ground state. Note that, in contrast to the dimensional reduction performed for number and energy damping[47], we do not require the thermal cloud to be three dimensional. The results in this section hence extend into the regime of deep traps. We can thus write in the two and one dimensional case
@@
The noises fulfill the correlations
@@
We derived a dimensionally reduced version of the evaporative damping. In contrast to the dimensional reduction performed for energy and number damping[47] this reduction stays valid in traps deeper than the thermal energy $k_{\text{B}}T$. As long as the thermal cloud stays three dimensional, number and energy damping include interactions with all the particles in the incoherent region[47], while evaporative damping only includes scattering with particles in the ground state along the reduced dimensions. However, the trapping implies an increase of the coherent region density relative to the incoherent region density. As more of the physics of evaporative damping happens in the coherent region compared to number and energy damping, in reduced dimensions, evaporative damping can still be expected to play a relevant role. For deep traps in which almost all particles sit in the ground state along the reduced dimension, it might actually be enhanced. In appendix F we estimate the strength of evaporative damping in one dimension at high phase space density and compare it with energy damping, indicating indeed a relevant role of evaporative damping in one dimension.
@@ VI Conclusions
A rough estimate of the evaporative damping strength close to the transition point (71) suggests that it is particularly strong compared to number damping for weakly interacting systems, for which the scattering length is much smaller than the de-Broglie wavelength. To verify the importance of evaporative damping in this regime will require future research.
In analyzing the importance of the evaporative damping process we focused on homogenous systems. We note that in traps the bulk of the coherent region and incoherent region become spatially seperated. The physics of the evaporative damping process is more strongly reliant on the coherent density and less affected by the density of the incoherent region compared to number and energy damping. Therefore, in trapped systems we can expect an increase in the relative strength of evaporative damping. Considering evaporative damping in a trap might hence be an important future direction of study.
We also derive a dimensionally reduced form of evaporative damping, equation (84). The reduction is valid if a deep harmonic trapping potential is applied along one (two) of the axes. It allows the study of the effect of evaporative damping on physics in reduced dimensions. One application could be the dynamics of stable non-linear excitations of the GPE, e.g. dark solitons [3, 4, 5, 48] in one dimension and vortices [42, 8, 9, 10] and Jones-Roberts solitons [6, 43, 7, 49] in two dimensions. We note that in dimensionally reduced systems the coherent density is increased relative to the thermal cloud. Accordingly, evaporative damping could dominate thermalization in dimensionally reduced systems, in consistency with an estimate for one dimensional systems at high phase space density in appendix F.
@@ Appendix E Linearization
Appendix F One Dimensional Reduced Dynamics at high Phase Space Density
In the one dimensional case we can calculate $\bar{\sigma}$ analogously to the three dimensional case outlined in IV.2. At high phase space density in a box we find
$\begin{split}\bar{\sigma}_{1}&=n_{C}\frac{2}{\pi}\frac{a_{\text{s}}^{2}\lambda_{\text{th}}^{2}N_{\text{cut}}}{k_{\text{cut}}l_{z}^{2}l_{y}^{2}}\int dv\cos(k_{\text{cut}}v)\\ &\times\langle\psi^{*}(x+v/2)\psi(x-v/2)\rangle\\ &\simeq\frac{2}{\pi}\frac{n_{C}a_{\text{s}}^{2}\lambda_{\text{th}}^{2}N_{\text{cut}}}{k_{\text{cut}}l_{z}^{2}l_{y}^{2}}\int dv\cos(k_{\text{cut}}v)\\ &\times n_{C}\exp\left(\left\langle\theta(x+v/2)\theta(x-v/2)-\theta(x)^{2}\right\rangle\right)\\ &\simeq\frac{2}{\pi}\frac{n_{C}^{2}a_{\text{s}}^{2}\lambda_{\text{th}}^{2}N_{\text{cut}}}{k_{\text{cut}}l_{z}^{2}l_{y}^{2}}\int dv\cos(k_{\text{cut}}v)\\ &\times\exp\left(\frac{1}{n_{C}\lambda_{\text{th}}^{2}}\int_{-k_{\text{cut}}}^{k_{\text{cut}}}dk\frac{\cos(kv)-1}{k^{2}}\right)\\ &=\frac{2}{\pi}\frac{n_{C}^{2}a_{\text{s}}^{2}\lambda_{\text{th}}^{2}N_{\text{cut}}}{k_{\text{cut}}^{2}l_{z}^{2}l_{y}^{2}}\int dy\cos(y)\\ &\times\exp\left(\frac{2}{n_{C}\lambda_{\text{th}}^{2}}\frac{1-\cos(y)-\text{Si}(y)y}{k_{\text{cut}}}\right)\end{split}$ (124)
For $n_{C}\lambda_{\text{th}}^{2}k_{\text{cut}}\ll 1$, as is the case at high phase space density and SPGPE validity, we can approximate the argument of the exponential by its asymptotic and find
$\begin{split}\bar{\sigma}_{1}&\simeq\frac{2}{\pi}\frac{n_{C}^{2}a_{\text{s}}^{2}\lambda_{\text{th}}^{2}N_{\text{cut}}}{k_{\text{cut}}^{2}l_{z}^{2}l_{y}^{2}}\int dy\cos(y)\exp\left(\frac{-\pi y}{n_{C}\lambda_{\text{th}}^{2}k_{\text{cut}}}\right)\\ &=4\frac{n_{C}a_{\text{s}}^{2}N_{\text{cut}}}{k_{\text{cut}}^{3}l_{z}^{2}l_{y}^{2}}.\end{split}$ (125)
For a symmetric trap $l_{y}=l_{z}$, we have $l_{z}^{2}=2n_{C}a_{\text{s}}\xi^{2}$ [47] and hence conclude
$\bar{\sigma}_{1}\simeq\frac{N_{\text{cut}}}{[k_{\text{cut}}\xi]^{3}n_{C}\xi}.$ (126)
In one dimension, energy damping can be cast into a local form [47] allowing a direct comparison to the energy damping like term in the (pseudo-)local form of evaporative damping (64), as the processes than only differ in the prefactor. The energy damping strength is given by
$\bar{\mathcal{M}}=2\sqrt{8\pi}\frac{a_{\text{s}}^{2}}{l_{z}}N_{\text{cut}}=4\sqrt{\pi}\frac{a_{\text{s}}^{2}N_{\text{cut}}}{\sqrt{n_{C}a_{\text{s}}}\xi},$ (127)
while the evaporative damping strength for the energy damping like term is determined by $4\bar{\sigma}_{1}/n_{C}$. The relative strength between energy and evaporative damping is hence given by
$\frac{n_{C}\bar{\mathcal{M}}}{4\bar{\sigma}_{1}}=\sqrt{\pi}[n_{C}a_{\text{s}}]^{3/2}[k_{\text{cut}}\xi]^{3}.$ (128)
With the ideal cut-off choice identified in [44] we can write this as
$\frac{n_{C}\bar{\mathcal{M}}}{4\bar{\sigma}_{1}}=\sqrt{\pi}[n_{C}a_{\text{s}}]^{3/2}\left[2\left(2\frac{k_{\text{B}}T}{\hbar\omega_{z}}\frac{1}{n_{C}a_{\text{s}}}\right)^{1/3}-2\right]^{3/2}.$ (129)
Energy damping maximizes compared to evaporative damping for
$n_{C}a_{\text{s}}=2(2/3)^{3}\frac{k_{\text{B}}T}{\hbar\omega_{z}}$ (130)
in which case it is
$\frac{n_{C}\bar{\mathcal{M}}}{4\bar{\sigma}_{1}}=2^{3/2}\left(\frac{2}{3}\right)^{9/2}\sqrt{\pi}\left(\frac{k_{\text{B}}T}{\hbar\omega_{z}}\right)^{3/2}\simeq 0.8\left(\frac{k_{\text{B}}T}{\hbar\omega_{z}}\right)^{3/2}.$ (131)
For the dimensional reduction to be valid, the thermal energy $k_{\text{B}}T$ should not be much larger than the transverse trap energy scale $\hbar\omega_{z}$. We conclude that in one dimensional systems, evaporative damping will often be at least comparable to energy damping and increases in importance with the depth of the transverse trap.
In two dimensions the slow algebraic decay of first order correlations[50] means that the derivation of the (pseudo-)local form becomes invalid. Hence, it hampers an analytic estimate of the strength of evaporative damping along the lines of the estimates in one and three dimensions.